R. Bachimanchi, S. Cooper, M. Drury, E. Daly, A. Freyberger, R. Geng, C. Hovater, L. King, G. Krafft, F. Marhauser, C. Mounts, T. Powers, T. Plawski, H. Wang, H. Zhang
Agenda and Notes
1. Operations Status and Summary
He Processing complete! Mike Drury will report 1 month from now.
Arne reports that preliminary assessment of Mike's data means that we should be able to run the C20s low enough that trip rates will be good.
RF soak status? Begun in the north. Cryo problems in the south. Ops drive highs have been increased to allow large trip rates so that the trip data for the newly He processes cavities may be acquired quickly. Slight SNAFU reported during the meeting. Radiation trips were stimulated by some of the high gradient running during the meeting.
2. Investigation team for heat and He pressure in R100 (Daly, Mounts)
Already follow-up from two weeks ago. L. King reports pressure transducers available now at either end of linacs and near "T"s in the middle of linac at He flow split. Presently for monitoring only. Maybe eventually could be used for active feedback.
3. C75 Project Summary Documentation (Krafft, Freyberger)
- Requirements Discussion
- First Draft version is attached below, modeled on tables in the
Cornell ERL "Project Definition Design Report". It contains input from Rimmer, Hutton, and Krafft but needs work on the cryomodule side. Daly and Mounts to help complete. Arne to derive cooling power requirement for charter document. Arne suggested publishing in tech note in about 6 weeks. Krafft will shepherd.
4. CEBAF Beamline Vacuum (Geng)
Complete Project Charter Document; Project Leader (Freyberg, Rimmer, Krafft, Hutton)
Data mining for field emission (Geng)
C75 parameter table! (krafft, rimmer, marhauser)
Heat For Project Charter (freyberger)